# Elucidated Rolling Diffusion Models for Probabilistic Weather Forecasting

### v1

@@ F.3 Comprehensive ERA5 power spectra
To assess the physical realism of the generated forecasts, we analyze their spectral properties, as presented in Fig. 14. This figure shows the spectral density of 14-day forecasts for several key atmospheric variables, averaged over high latitudes ($[60^{\circ},90^{\circ}]$). The absolute spectra (subfigure (a)) and spectra normalized by ERA5 reanalysis (subfigure (b)) consistently demonstrate the high fidelity of ERDM. Our model’s spectra closely align with the ERA5 target across a broad range of zonal wavenumbers and frequently match or even slightly outperform those from the operational physics-based IFS ENS model. This accurate representation indicates that ERDM effectively captures the energy distribution from large planetary scales down to smaller synoptic scales for diverse atmospheric fields. In contrast, NeuralGCM ENS exhibits a marked underestimation of energy at mid to high frequencies for several variables, evident by its spectra decaying more rapidly in the absolute plots and falling below unity in the normalized plots at these scales. The robust spectral characteristics of ERDM across multiple variables underscore its capability to produce physically consistent and realistic long-range weather forecasts.
@@
Figure 14: Spectral density of 14-day forecasts, averaged over high latitudes ($[60^{\circ},90^{\circ}]$). ERDM generates highly accurate spectra that match or even slightly beat the physics-based model IFS ENS, while NeuralGCM ENS underestimates energy at mid to high frequencies. Subfigure (a) displays the absolute spectral density, and subfigure (b) presents the normalized spectral density (absolute spectra divided by the target ERA5 spectra).
@@ F.4 ERA5 forecast visualizations
See Fig. 15 and Fig. 16 for visualizations of example forecasts by ERDM for a specific initial condition (2022-01-01) and several lead times. As suggested by the good power spectra (see section above), ERDM produces realistic forecasts, including at long lead times (e.g., 15 days).

### v3

@@ F.3 Comprehensive ERA5 power spectra
To assess the physical realism of the generated forecasts, we analyze their spectral properties, as presented in Fig.˜14. This figure shows the spectral density of 14-day forecasts for several key atmospheric variables, averaged over high latitudes ($60^{\circ},90^{\circ}]$). The absolute spectra (subfigure (a)) and spectra normalized by ERA5 reanalysis (subfigure (b)) consistently demonstrate the high fidelity of ERDM. Our model’s spectra closely align with the ERA5 target across a broad range of zonal wavenumbers and frequently match or even slightly outperform those from the operational physics-based IFS ENS model. This accurate representation indicates that ERDM effectively captures the energy distribution from large planetary scales down to smaller synoptic scales for diverse atmospheric fields. In contrast, NeuralGCM ENS exhibits a marked underestimation of energy at mid to high frequencies for several variables, evident by its spectra decaying more rapidly in the absolute plots and falling below unity in the normalized plots at these scales. The robust spectral characteristics of ERDM across multiple variables underscore its capability to produce physically consistent and realistic long-range weather forecasts.
@@
Figure 14: Spectral density of 14-day forecasts, averaged over high latitudes ($60^{\circ},90^{\circ}]$). ERDM generates highly accurate spectra that match or even slightly beat the physics-based model IFS ENS, while NeuralGCM ENS underestimates energy at mid to high frequencies. Subfigure (a) displays the absolute spectral density, and subfigure (b) presents the normalized spectral density (absolute spectra divided by the target ERA5 spectra).
@@ F.4 ERA5 forecast visualizations
See Fig.˜15 and Fig.˜16 for visualizations of example forecasts by ERDM for a specific initial condition (2022-01-01) and several lead times. As suggested by the good power spectra (see section above), ERDM produces realistic forecasts, including at long lead times (e.g., 15 days).